# Search for a gravitational wave background from primordial black hole binaries using data from the first three LIGO-Virgo-KAGRA observing runs

### v1

@@ Abstract
Introduction.— Gravitational-wave (GW) astronomy has the potential to revolutionize our understanding of the Universe by providing insights into its most energetic and dynamic phenomena. GW detections allow us to explore regions of space-time that are inaccessible through electromagnetic observations. Since the first binary black hole (BBH) merger detected by the LIGO-Virgo-KAGRA (LVK) collaboration [1], the field has rapidly expanded with numerous GW events reported over the past decade [2, 3, 4]. These observations have advanced our knowledge of compact objects, their merger mechanisms, and have enabled tests of general relativity in the strong-field regime [5]. Additionally, GW signals have been used to measure the Hubble constant through standard sirens [6].
@@
Several previous studies have investigated the possibility that some of the detected CBC events originate from PBH mergers [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 35, 41, 42, 43]. These studies typically analyze the observed population of CBCs obtained through individual detections by examining their mass distributions, spin orientations, and redshift distributions to discuss whether the properties of these events are consistent with an astrophysical origin or suggest a primordial origin. For example, a BBH merger with component masses in the so-called “mass gap” region—where standard stellar evolution models predict a scarcity of black holes due to pair-instability supernovae—has prompted speculation about a primordial origin [44, 32, 39, 32, 45, 31]. Furthermore, a subsolar mass event, can serve as distinct evidence of PBHs [46, 47, 48].
@@
Another approach is to search unresolvable binaries as a GWB. Some studies have explored constraints on PBHs by considering the GWB produced by their mergers, a line of research pioneered in [50]. Constraints have been discussed in the literature [26, 28, 51, 52, 53, 54, 55], involving typically a rough comparison between the theoretical predictions and the upper bounds obtained from the GWB search performed by the LVK collaboration [56, 57, 58], which assumes a power-law spectrum $\Omega_{\rm GW}\propto f^{\alpha}$. However, for GWB generated by compact binaries with mass distributions extending beyond $\sim 50M_{\odot}$, the spectrum deviates from a simple power law.
@@
Motivated by this, we analyze data from the first three LVK observing runs (O1–O3) [56, 57, 58] and perform Bayesian parameter inference for the GWB from PBHs with a wide range of masses, including sub-solar masses up to $10^{3}$ solar masses. This allows us to account for the complete spectral shape and explore a broad parameter space that extends beyond the detection of individual mergers by LVK. A recent work in this direction is [43], although it is primarily focused on distinguishing between astrophysical and primordial BBHs. In addition, the PBH model used in [43] was limited to a fixed mass of $1$ and $30M_{\odot}$. It is also important to note that, for the first time, we consider in our study the combination of early and late binary formation channels to get corresponding constraints.
@@
In what follows, we aim to effectively bridge the gap between theoretical studies and observational data through robust Bayesian analysis. Using the most up-to-date publicly available data from the LVK (O1-O3) searches, we provide more reliable and comprehensive constraints on PBH abundance for peaked mass distributions across a broad mass range, a topic gaining significant attention in the field, particularly in relation to other cosmological constraints.
@@
Modeling.— To parameterize the abundance of PBH, we commonly use $f_{\rm PBH}=\Omega_{\rm PBH}/\Omega_{\rm DM}$, which is the fraction of PBH in the dark-matter energy density today. One of the key factors influencing the GW spectrum in our model is the mass function of PBHs. It is commonly parametrized using monochromatic or Log-Normal distributions. They may not accommodate a mass function with a detailed shape, such as the one associated with the QCD phase transition [59, 60, 61], but it is a good approximation for PBHs arising from a peak in the primordial power spectrum [62].
Using the PBH energy density $\rho_{\rm PBH}$, we define the Log-Normal PBH mass function $p(m)$ as
@@
The mass function is normalized to satisfy $\int p(m)\mathrm{d}\ln m=1$.
@@
Early binary formation.— Early binaries are formed in the radiation-dominated era shortly after PBH formation. In this scenario, a binary is formed from a pair of closely spaced PBHs with the surrounding third object acting on the pair via a tidal force that generates the angular momentum of the binary [65, 21]. The merging rate (per unit logarithmic mass of the two binary components) of early binaries at a time $t$ is given by [66, 27, 24, 21, 67].
@@
where $m_{1}$ and $m_{2}$ are the two black hole masses and $t_{0}$ is the age of the Universe. A rate suppression factor, $S$, has been introduced to account for two principal mechanisms that impede binary formation due to local matter inhomogeneities and nearby PBHs $S_{1}(m_{1},m_{2},f_{\rm PBH})$ and the clustering due to their initial Poissonian fluctuations $S_{2}(f_{\rm PBH})$ [27, 28].
Analytical prescriptions for these suppression factors have been studied in [28, 27]. We present these expressions in the Appendix A, including the ranges relevant to our work. Notably, for narrow mass distributions with $\sigma<1$, the suppression factor $S$ is less than 0.2 (see Fig. 2). In this study, we fix $S$ as $S=S_{1}\times S_{2}=0.1$. This choice is reasonable in the range $0.010.001<f_{\rm PBH}<$ 0.01, where our constraints are most applicable. We have confirmed that our specific choice of $S$ does not significantly affect our results. As detailed in the Appendix A, the suppression factor has significant uncertainties, particularly in the case of broad mass distributions, which led us to focus on the scenario of a narrow mass distribution in this paper.
@@
Late binary formation.— Late binaries form much later due to dynamical capture in the clusters of PBHs during the matter-dominated era [20]. For a general mass distribution of PBHs, an expression for the merger rate has been proposed in the literature [68, 69]:
$\frac{d^{2}R_{\rm LB}}{d\ln m_{1}d\ln m_{2}}=\frac{R_{\rm clust}}{\rm Gpc^{3}yr}f_{\rm PBH}^{2}\frac{(m_{1}+m_{2})^{10/7}}{(m_{1}m_{2})^{5/7}}\,,$ (6)
where $R_{\rm clust}=400$ (see the discussion in [31] for this choice) is a scaling factor that encapsulates the clustering properties and velocity distribution of the PBHs. This relationship is derived from the two-body capture process occurring within a cluster and operates under the assumption that the time required for the binary system to merge is significantly shorter than the age of the universe.
Large uncertainties remain on the exact value of $R_{\rm clust}$ and its possible mass and PBH model dependence. In our case, the rate of late binaries remains subdominant compared to early binaries, except for $m_{\rm PBH}\gtrsim 200M_{\odot}$. Using GWB searches to set new limits on the PBH abundance for a large mass range is an interesting avenue that one could follow using the methodology we present here. One can note that the limits obtained on $f_{\rm PBH}$ using our benchmark value of $R_{\rm clust}$ can be redrawn for any other value, with the simple rescaling $f_{\rm PBH}\longleftrightarrow f_{\rm PBH}(400/R_{\rm clust})^{1/2}$. This way, the limits we obtain can in fact become much more stringent in the case of smaller and denser PBH clusters.
@@
Methodology.— Let us outline the Bayesian framework employed in our search for GWB originating from PBHs and other compact binary sources. Our methodology was first introduced in [70] and closely follows that of previous studies [71, 72, 73, 74, 75, 76, 77, 78, 79, 57, 58], with specific adaptations to our analysis.
@@
We utilize data from the Advanced LIGO detectors located in Hanford (H) and Livingston (L), as well as data from the Virgo (V) observatory, during the first three observing runs (O1–O3) [80, 81, 82, 83]. Our search incorporates all available baselines (HL, HV, LV) (for O1 and O2 we use only HL) to maximize sensitivity and leverage the complementary geographical locations of the detectors.
@@
Prior Distributions.— We employ two distinct sets of priors in our analysis. The first set, referred to as Agnostic Log-Normal, includes the parameters: the mean of the Log-Normal mass distribution $\mu$, its width $\sigma$, and the PBH fraction $f_{\rm PBH}$ in the dark matter energy density today, as introduced previously. The second set of priors, termed Mono-chromatic (MN), is identical to the Agnostic Log-Normal (ALN) priors except that the width $\sigma$ is fixed at $10^{-2}\,M_{\odot}$.
@@
Figure 1: Two-sigma ($2\sigma$) constraints on the PBH fraction ($f_{\rm PBH}$) as a function of $M_{\rm PBH}$($=\mu$) from the search for the gravitational-wave background (GWB) using the ALN (MN) constraint is shown in solid (dashed) red. We also display supernova lensing constraints (SNe) [86] in purple, Massive Compact Halo Object (MACHO) constraints [87] in orange designated as (EROS), the recent Optical Gravitational Lensing Experiment (OGLE) constraints are shown in green with different Milky Way (MW) models [88], and cosmic microwave background (CMB) constraints from [89]. We used the PBHBounds [90] to generate this plot.
@@
For both sets we also take into account for an additional CBC background from astrophysical black holes. Given their typical mass range of ${\cal O}(1-10)M_{\odot}$, the spectrum is well approximated by a simple power law characterized by a single parameter, and is expressed as $\Omega_{\rm cbc}$ as $\Omega_{\rm CBC}(f)=\Omega_{\rm cbc}(f/f_{\rm ref})^{2/3}$, with $f_{\rm ref}=25$ Hz [91]. Therefore, our model is $\Omega_{\rm GW}=\Omega_{\rm CBC}+\Omega_{\rm EB}+\Omega_{\rm LB}$.
The prior distributions for both sets of model parameters are specified in Table  1. These ranges encompass plausible values based on theoretical models and existing constraints [26, 91, 28]. Note that we focus on relatively sharp mass distributions due to uncertainties in the theoretical predictions for broad mass distributions as already mentioned. Thus, the maximum value of $\sigma$ for ALN prior is set to 1.
@@
Bayesian Search.— In order to perform Bayesian parameter inference, we employ the nested sampling algorithm from the dynesty package [92] for robust parameter space exploration. To expedite the computation of integrals in the gravitational wave spectrum calculations, we utilize Sobol Monte Carlo integration [93], a quasi-random low-discrepancy sequence method that improves convergence rates over traditional Monte Carlo approaches, thereby reducing computational overhead. To account for calibration uncertainties in GW detector data, we use the Python library pygwb [75], which integrates these uncertainties into the parameter estimation process.
Parameter ALN Prior MN Prior
$\Omega_{\rm cbc}$ $\mathcal{LU}$($10^{-10}$, $10^{-7}$) $\mathcal{LU}$($10^{-10}$, $10^{-7}$)
$\mu\,(M_{\odot})$ $\mathcal{LU}$($10^{-1}$, $10^{3}$) $\mathcal{LU}$($10^{-1}$, $10^{3}$)
$\sigma$ $\mathcal{LU}$($10^{-2}$, $10^{0}$) Fixed at $10^{-2}$
$f_{\rm PBH}$ $\mathcal{LU}$($10^{-5}$, $1$) $\mathcal{LU}$($10^{-5}$, $1$)
@@
Table 1: Prior distributions for the model parameters in the ALN and MN frameworks used in the Bayesian analysis. $\mathcal{LU}$ denotes a LogUniform distribution. For the MN framework, $\sigma$ is fixed at $10^{-2}$.
@@
The results indicate that no significant GWB associated with PBHs or CBCs has been detected over the noise hyphothesis, with logarithmic Bayes factors $\sim-1$. We establish upper limits on the CBC energy density parameter, $\Omega_{\rm cbc}$, of $6.0\times 10^{-9}$ for ALN priors and $6.2\times 10^{-9}$ for MN priors at the $2\sigma$ level which are consistent with LVK contraints [58].
From the non-detection, we can constrain the fraction of PBHs, $f_{\rm PBH}$, as a function of the mean mass $\mu$. The obtained constraint, along with other studies from lensing surveys like EROS ([87]), OGLE ([88]), supernova Ia ([86]) and cosmic microwave background (CMB) constraints are summarized in Fig. 1. We also provide the resulting corner plots for all model parameters in Appendix C.
The constraint on $f_{\rm PBH}$ is obtained by marginalizing over the mass function width $\sigma$ in the ALN scenario. However, because of the relatively narrow range of $\sigma$, our analysis shows no significant differences between the two distinct prior sets when imposing constraints on $f_{\rm PBH}$.
These findings demonstrate that GWB constraints are and will continue to be a vital probe, particularly within the mass range $[10^{1},3\times 10^{2}]\,M_{\odot}$. This particular mass range is dominated by the early binaries and our constraints remain between $0.010.001<f_{\rm PBH}<$ 0.01. For $M_{\rm PBH}>10^{2}M_{\odot}$, the GWB spectral amplitude is dominated by the late binary formation channel. Although the LVK loses its constraining power sharply in this mass range due to the limited frequency range, this suggests an interesting possibility for constraining this mass range using the late binary formation channel with the Einstein Telescope [94, 95] and Cosmic Explorer[96, 97], which have greater sensitivity at low frequencies.
Conclusion.— In this study, we conducted a targeted search for a GWB originating from PBHs using cross-correlation data from the first three observing runs of the LVK Collaboration [56, 57, 58]. Unlike previous research that focused on the population of resolved gravitational-wave events, we utilized cross-correlated LVK data specifically to search for a GWB resulting from the superposition of many unresolved binary events. Our mass range encompasses PBH binaries with component masses both below and above those detected in the LVK catalog [2, 4].
@@
Acknowledgments.— The authors would like to thank Vuk Mandic and the LVK Stochastic group for comments.

### v2

@@ Abstract
.—I Introduction
Gravitational-wave (GW) astronomy has the potential to revolutionize our understanding of the Universe by providing insights into its most energetic and dynamic phenomena. GW detections allow us to explore regions of space-time that are inaccessible through electromagnetic observations. Since the first binary black hole (BBH) merger detected by the LIGO-Virgo-KAGRA (LVK) collaboration [1], the field has rapidly expanded with numerous GW events reported over the past decade [2, 3, 4]. These observations have advanced our knowledge of compact objects, their merger mechanisms, and have enabled tests of general relativity in the strong-field regime [5]. Additionally, GW signals have been used to measure the Hubble constant through standard sirens [6].
@@
Several previous studies have investigated the possibility that some of the detected CBC events originate from PBH mergers [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 35, 41, 42, 43]. These studies typically analyze the observed population of CBCs obtained through individual detections by examining their mass distributions, spin orientations, and redshift distributions to discuss whether the properties of these events are consistent with an astrophysical origin or suggest a primordial origin. For example, a BBH merger with component masses in the so-called “mass gap” region—where standard stellar evolution models predict a scarcity of black holes due to pair-instability supernovae—has prompted speculation about a primordial origin [44, 32, 39, 32, 45, 31]. Furthermore, a subsolar mass event, can serve as distinct evidence of PBHs [46, 47, 48, 49].
@@
Another approach is to search unresolvable binaries as a GWB. Some studies have explored constraints on PBHs by considering the GWB produced by their mergers, a line of research pioneered in [51]. Constraints have been discussed in the literature [26, 28, 52, 53, 54, 55, 56, 57], typically involving a rough comparison between theoretical predictions and the upper bounds obtained from the isotropic GWB search performed by the LVK collaboration [58, 59, 60], which assumes a power-law spectrum $\Omega_{\rm GW}\propto f^{\alpha}$. Although this assumption may be valid for a GWB originating from an unresolved BBH population from Population I or II stars (where lower end is bounded by the Chandrasekhar limit, while the intermediate mass range is constrained by pair-instability), PBH mass distributions can extend significantly beyond $\sim 50\,M_{\odot}$. As a result, the GWB spectrum could deviate from a simple power-law form when PBH sources are considered ([57]).
@@
Motivated by these, we analyze public GWB data from the first three LVK observing runs (O1–O3) [58, 59, 60] and perform a full Bayesian parameter inference for the GWB from PBHs with a wide range of masses, including sub-solar masses up to $10^{3}$ solar masses. This allows us to account for the complete spectral shape and explore a broad parameter space that extends beyond the detection of individual mergers by LVK. Some recent works in this direction are [43, 57], although the former was primarily focused on distinguishing between astrophysical and primordial BBHs. In addition, the PBH model used in [43, 57] were limited to a single mass bin of $1$ and $30M_{\odot}$. It is worth noting that in what follows it is for the first time considered the combination of different formation channels to get corresponding constraints and assess their relative significance.
@@
In what follows, we aim to effectively bridge the gap between theoretical studies and observational data through robust Bayesian analysis. Updating to the most up-to-date publicly available data from the LVK (O1-O3) searches, we provide more reliable and comprehensive constraints on PBH abundance for peaked mass distributions across a broad mass range, a topic gaining significant attention in the field, particularly in relation to other cosmological constraints. The Bayesian framework offers key advantages in this context. It is a natural framework to include various constraints between parameters an uncertainties on them into the analysis. Moreover, in the event of a future GWB detection, Bayesian inference will be essential for identifying the underlying source population and distinguishing between astrophysical and primordial origins.
The paper is organized as follows. In Sec. II, we describe the models for calculation of the PBH merger rate. In Sec. III, we outline our Bayesian inference method and present the results. Then we conclude in Sec. IV. Appendix A provides additional details on the suppression factors relevant to the early binary formation scenario. AppendixB describes the cross-correlation statistics used in our analysis. Appendix C presents an injection study exploring various potential scenarios. Appendix D provides the posterior distributions of all parameters corresponding to the main results.
.—II Modeling
@@
To parameterize the abundance of PBH, we commonly use $f_{\rm PBH}=\Omega_{\rm PBH}/\Omega_{\rm DM}$, which is the fraction of PBH in the dark-matter energy density today. One of the key factors influencing the GW spectrum in our model is the mass function of PBHs. It is commonly parametrized using monochromatic or Log-Normal distributions. They may not accommodate a mass function with a detailed shape, such as the one associated with the QCD phase transition [61, 62, 63], but it is a good approximation for PBHs arising from a peak in the primordial power spectrum [64].
Using the PBH energy density $\rho_{\rm PBH}$, we define the PBH mass function $p(m)$ as
@@
The mass function is normalized to satisfy $\int p(m)\mathrm{d}\ln m=1$ and is usually motivated by inflationary scenarios [65] . We concentrate on relatively narrow mass distributions ($\sigma<1$) because theoretical predictions for broader distributions contain significant uncertainties [66, 64, 67, 68]. Nevertheless, our chosen distribution remains sufficiently broad within the LVK sensitivity range, allowing mass variations on the order of the central value ($\sim\mathcal{O}(\mu)$), potentially spanning several tens of solar masses around $\sim 50\,M_{\odot}$. While our simple log-normal distribution represents progress towards understanding these mass distributions, detailed modeling incorporating more complex structure will be deferred to future work.
@@
II.1 Early binary formation
Early binaries are formed in the radiation-dominated era shortly after PBH formation. In this scenario, a binary is formed from a pair of closely spaced PBHs with the surrounding third object acting on the pair via a tidal force that generates the angular momentum of the binary [71, 21]. The merging rate (per unit logarithmic mass of the two binary components) of early binaries at a time $t$ is given by [72, 27, 24, 21, 73].
@@ II.1 Early binary formation
where $m_{1}$ and $m_{2}$ are the two black hole masses and $t_{0}$ is the age of the Universe. A rate suppression factor, $S$, has been introduced to account for two principal mechanisms that impede binary formation due to local matter inhomogeneities and nearby PBHs $S_{1}(m_{1},m_{2},f_{\rm PBH})$ and the clustering due to their initial Poissonian fluctuations $S_{2}(f_{\rm PBH})$ [27, 28]. Analytical prescriptions for these suppression factors have been studied in [28, 27] which are shown to be consistent with numerical simulations [27]. We detail these aspects in Appendix A, including the relevant ranges for our work and their limitations.
II.2 Late binary formation
@@
Late binaries form much later due to dynamical capture in the clusters of PBHs during the matter-dominated era [20]. For a general mass distribution of PBHs, an expression for the merger rate has been proposed in the literature [74, 75]:
$\frac{d^{2}R_{\rm LB}}{d\ln m_{1}d\ln m_{2}}=\frac{R_{\rm clust}}{\rm Gpc^{3}yr}f_{\rm PBH}^{2}\frac{(m_{1}+m_{2})^{10/7}}{(m_{1}m_{2})^{5/7}}\,.$ (6)
$R_{\rm clust}$ is a single parameter that encapsulates how much the local environment of PBHs boosts their merger rate which depends on the velocity dispersion and density contrast in PBH. In $\Lambda\rm CDM$ model without significant small-scale Poisson fluctuations, one typically finds low values in the range of a $O(1\sim 10)$ [76] by applying the Press-Schechter formalism to the linear matter power spectrum. However, if PBHs make up most of the dark matter at around solar masses or above, their discrete (Poisson-like) distribution can seed dense subhalos that raise this factor into the $O(10^{2}\sim 10^{3})$ [31, 77]. Current GW data from individually detected binaries, particularly from mergers near the neutron-star and pair-instability mass gaps, is consistent with $\sim 4\times 10^{2}$ [31]. This value remains somewhat model-dependent, since processes like direct cluster formation or varying PBH mass fractions can push it higher or lower, but the data suggest that a few hundred is a reasonable benchmark. To encapsulate these different cases, we will take three reference values for $R_{\rm clust}=[1,4\times 10^{2},10^{3}]$.
.—III Bayesian search
III.1 Methodology
@@
Let us outline the Bayesian framework employed in our search for GWB originating from PBHs and other compact binary sources. Our methodology was first introduced in [78] and closely follows that of previous studies [79, 80, 81, 82, 83, 84, 85, 86, 87, 59, 60], with specific adaptations to our analysis.
@@
We utilize data from the Advanced LIGO detectors located in Hanford (H) and Livingston (L), as well as data from the Virgo (V) observatory, during the first three observing runs (O1–O3) [88, 89, 90, 91]. Our search incorporates all available baselines (HL, HV, LV) (only HL baseline is available for O1 and O2) to maximize sensitivity and leverage the complementary geographical locations of the detectors.
@@ III.2 Prior Distributions
We employ a set of priors includes the parameters: the mean of the Log-Normal mass distribution $\mu$, its width $\sigma$, and the PBH fraction $f_{\rm PBH}$ in the dark matter energy density today, as introduced previously.
@@
Figure 1: (Left) 95% C.L. constraints on $f_{\rm PBH}$ as a function of $M_{\rm PBH}$($=\mu$) from the search for the GWB using the different $R_{\rm clust}$ choices are shown in red ($1$-dashed, $4\times 10^{2}$-solid, $10^{3}$-dotted). We also display supernova lensing constraints (SNe) [94] in purple, Massive Compact Halo Object (MACHO) constraints [95] in orange designated as (EROS), the recent Optical Gravitational Lensing Experiment (OGLE) constraints are shown in green with different Milky Way (MW) models [96], and cosmic microwave background (CMB) constraints from [97]. We used the PBHBounds [98] to generate this plot. (Right) Zoomed‑in view of the same plot.
@@
We also take into account for an additional otherwise individually unresolvable CBC background from astrophysical black holes which can include lighter binaries like neutron stars as well. Their spectrum is well approximated by a simple fiducial power law characterized by a single parameter, and is expressed as $\Omega_{\rm cbc}$ as $\Omega_{\rm CBC}(f)=\Omega_{\rm cbc}(f/f_{\rm ref})^{2/3}$, with $f_{\rm ref}=25$ Hz [99, 58, 59, 60]. Deviations from this fiducial power law are not prominent and the expected CBC spectrum is only deviates from the power law beyond the current sensitivity (e.g. Fig. 5 in [60]). Therefore, our model is $\Omega_{\rm GW}=\Omega_{\rm CBC}+\Omega_{\rm EB}+\Omega_{\rm LB}$.
The prior distributions for the model parameters are specified in Table  I. These ranges encompass plausible values based on theoretical models and existing constraints [26, 28, 60, 31, 77, 35, 38, 42, 45]. We adopt a log‐uniform prior for the PBH abundance $f_{\rm PBH}$ because a Jeffreys (log‐uniform) prior is the least informative choice for a scale parameter—it assigns equal weight to each decade avoids biasing the inference toward any particular order of magnitude [100]. We set the lower bound on $f_{\rm PBH}$ prior to $10^{-5}$ which is usually the standard choice [45, 101], this makes our results directly comparable to the existing literature.
III.3 Result
@@
In order to perform Bayesian parameter inference, we employ the nested sampling algorithm from the dynesty package [102] for robust parameter space exploration. To expedite the computation of integrals in the GW spectrum calculations, we utilize Sobol Monte Carlo integration [103], a quasi-random low-discrepancy sequence method that improves convergence rates over traditional Monte Carlo approaches, thereby reducing computational overhead. To account for calibration uncertainties in GW detector data, we use the Python library pygwb [83], which integrates these uncertainties into the parameter estimation process.
Parameter Prior
$\Omega_{\rm CBC}$ $\mathcal{LU}$($10^{-10}$, $10^{-7}$)
$\mu\,(M_{\odot})$ $\mathcal{LU}$($10^{-1}$, $10^{3}$)
$\sigma$ $\mathcal{LU}$($10^{-2}$, $10^{0}$)
$f_{\rm PBH}$ $\mathcal{LU}$($10^{-5}$, $1$)
@@
Table I: Prior distributions for the model parameters in in the Bayesian analysis. $\mathcal{LU}$ denotes a LogUniform distribution.
@@
The results indicate that no significant GWB associated with PBHs or CBCs has been detected over the noise hypothesis, with logarithmic Bayes factors $\sim-1$. We establish upper limits on the CBC energy density parameter, $\Omega_{\rm cbc}$, of $\sim 3\times 10^{-9}$ at the $95\%$ C.L. level which are consistent with LVK constraints [60].
From the non-detection, we can constrain the fraction of PBHs, $f_{\rm PBH}$, as a function of the mean mass $\mu$. The obtained constraint, along with other studies from lensing surveys like EROS ([95]), OGLE ([96]), supernova Ia ([94]) and cosmic microwave background (CMB) constraints for different $R_{\rm clust}$ choices are summarized in Fig. 1 and II. We also provide the resulting corner plots for all model parameters in Appendix D.
The constraint on $f_{\rm PBH}$ is obtained by marginalizing over the mass function width $\sigma$ and $\Omega_{\rm CBC}$.
For $M_{\rm PBH}\gtrsim 3\times 10^{2}M_{\odot}$, the GWB spectral amplitude is dominated by the late binary formation channel. Although the LVK loses its constraining power sharply in this mass range due to the limited frequency range, this suggests an interesting possibility for constraining this mass range using the late binary formation channel with the Einstein Telescope [104, 105] and Cosmic Explorer[106, 107], which have greater sensitivity at low frequencies. We also demonstrate that Advanced LIGO A+ can provide valuable information about the underlying background. See Appendix C for one such scenario.
These findings demonstrate that GWB constraints are and will continue to be a vital probe, particularly within the mass range $[10^{1},3\times 10^{2}]\,M_{\odot}$. Our results improve the previous GWB constraints [28, 41, 52] by $O(2\sim 10)$ at $100\,M_{\odot}$ with the latest publicly available LVK observing run data.
Table II: 95% C.L. on $f_{\rm PBH}$ at various masses.
$R_{\rm clust}$ $\mu$ $1\,[M_{\odot}]$ $30\,[M_{\odot}]$ $10^{3}\,[M_{\odot}]$
$1$ $3.6\times 10^{-2}$ $5.3\times 10^{-3}$ $1.0\times 10^{0}$
$4\times 10^{2}$ $4.0\times 10^{-2}$ $6.2\times 10^{-3}$ $5.8\times 10^{-1}$
$10^{3}$ $3.3\times 10^{-2}$ $5.8\times 10^{-3}$ $4.8\times 10^{-1}$
.—IV Conclusion
In this study, we conducted a targeted search for a GWB originating from PBHs using cross-correlation data from the first three observing runs of the LVK Collaboration [58, 59, 60]. Unlike previous research that focused on the population of resolved GW events, we utilized cross-correlated LVK data specifically to search for a GWB resulting from the superposition of many unresolved binary events. Our mass range encompasses PBH binaries with component masses both below and above those detected in the LVK catalog [2, 4].
@@
Acknowledgments
The authors would like to thank Vuk Mandic and the LVK Stochastic group for comments.